# Fact-Checking the Output of Large Language Models via Token-Level Uncertainty Quantification

### v1

@@ 5 Benchmark for Evaluation of Claim-Level UQ Methods
We evaluate claim-level UQ techniques and their ability to spot hallucinations on the task of generating biographies. To make the evaluation as close as possible to the real-world scenario, we allow unrestricted generation of biographies from LLMs, sampling multiple outputs from a given LLM. This complicates the automatic evaluation of the fact-checking pipeline, because obtaining gold standard annotation require that we manually annotate all outputs from each model. Therefore, in addition to manual annotation, we annotate claims in generated texts automatically using the FactScore fact-checking tool Min et al. (2023), which has access to an external knowledge source. Such an approach enables completely automatic evaluation and allows us to scale up our experiments.
We generate LLM responses in English, Chinese, and Arabic to $100$ biography prompts. The typical biography prompt is Give me a biography for <person name> in different languages. The set of people was generated by asking GPT-4 to list the most famous people since 1900. The maximum generation length is set to $256$ tokens. If the last sentence of the generation is unfinished (i.e. does not end with any punctuation), it is discarded. We generate responses for the following LLMs: (for English) Vicuna 13b Zheng et al. (2023), Mistral 7b Jiang et al. (2023), Jais 13b Sengupta et al. (2023), and GPT-3.5-turbo Ouyang et al. (2022); (for Chinese) Yi 6B Yi (2023); (for Arabic) Jais 13b and GPT-4.
@@
We then decompose the generated text into atomic claims using ChatGPT. For each atomic statement, we map all its words back to generated text to access the corresponding token logits. Not all claims perfectly match to the original response. For example, for Vicuna 13b around $5\%$ of all claims do not successfully match because ChatGPT abstained to respond or outputted words not present in the original. We consider only successfully matched claims.
For English, we are able to perform annotation completely automatically: each atomic claim is classified by FactScore as supported or not supported. The underlying FactScore model is “retrieval+ChatGPT” with a dump of Wikipedia articles as an external knowledge source. We also manually annotated 100 English biographies produced by the Vicuna 13b model, 183 claims from Arabic biographies, and 1,603 claims in Chinese. Each of the statements was checked by two annotators with access to the corresponding Wikipedia article. The final label is set to ‘supported’ only if both annotators label it as so.
The statistics of the resulting datasets with automatically-labeled claims in English are presented in Table 4, and the statistics of manually annotated datasets is presented in Table˜5. As we can see, the majority of claims in the model output are correct, with 6-29% of hallucinations.
The automatic dataset construction process using FactScore is illustrated in Figure 6 in the Appendix.
@@ 6.1 Experimental Setup
Fact-checking of atomic claims is framed as a binary classification task, where uncertainty scores serve as predictors of non-factuality, and FactScore or human labels serve as ground truth. The evaluation metric is ROC-AUC.
For the CCP method, NLI scores are calculated using the DeBERTa-large model He et al. (2021) fine-tuned for this task. The number of alternatives used in CCP is $K=10$, except for GPT-3.5-turbo and GPT-4, as the OpenAI API does not allow us to retrieve more than $5$ alternatives from the token distribution. Details on expenses are presented in Appendix H.
@@ 6.2 Results for English on the FactScore Annotation
The main results of the experiments with FactScore labels for English are presented in Table 1. The proposed CCP method outperforms all other UQ techniques for each of the considered LLMs. The best overall result is obtained for Jais 13b, where CCP outperforms the closest competitor by 0.07 ROC-AUC.
We further analyze the performance by plotting it as a function of the number of considered sentences from the beginning of the generated text (see Figure 3). We note that the quality of each method decreases as the number of considered sentences increases. This may be due to the fact that the model tends to start the response with easy-to-know and hence reliable claims and as it generates more text, it has to produce more complex and less reliable statements, which is illustrated for the Vicuna 13b model in Figure 8 in the Appendix. For the majority of cases, CCP outperforms other methods except when we consider only the first two and the first five sentences generated by GPT-3.5-turbo.
@@ 6.3 Multilingual Results on Manual Annotation
Multilingual results based on manual annotation are presented in Tables˜2 and 3 and Figure˜9 in the Appendix.
Using manual annotation for English, we can also evaluate the performance of FactScore itself. The accuracy of automatic annotation is $77.2\%$ and ROC-AUC is $0.72$. A detailed analysis of FactScore mistakes is presented in Appendix F. For human annotation, the performance of all UQ methods appears to be even slightly higher than for the labels obtained using FactScore (Table 2) on the same set. Moreover, CCP outperforms FactScore itself by 0.06 ROC AUC. These results demonstrate that in the task of detecting LLM hallucinations, UQ techniques can be a strong alternative to fact-checking tools with an external knowledge source.
For Chinese and Arabic, CCP also outperforms the baselines. For the Chinese Yi 6b model, from Figure 9, we can see that the gap between CCP and the baselines is especially big for the several first claims. When considering more claims in the LLM output, CCP still clearly outperforms the maximum probability baseline, but the P(True) baseline substantially reduces the gap. For Jais and Arabic, CCP outperforms the closest competitor by 0.05 ROC AUC. On GPT-4 generations for Arabic, the metrics for all methods are low. We explain this observation by a small ratio of non-factual claims in the GPT-4 output. Compared to the other models, GPT-4 hallucinates much less.
@@
Model
Yi 6b,
Chinese
Jais 13b,
Arabic
GPT-4,
Arabic
CCP (ours) 0.64 $\pm$ 0.03 0.66 $\pm$ 0.02 0.56 $\pm$ 0.05
Maximum Prob. 0.52 $\pm$ 0.03 0.59 $\pm$ 0.02 0.55 $\pm$ 0.08
Token Entropy 0.57 $\pm$ 0.05 0.61 $\pm$ 0.02 0.48 $\pm$ 0.06
P(True) 0.63 $\pm$ 0.04 0.61 $\pm$ 0.02 0.50 $\pm$ 0.06
@@ 6.4 Ablation Studies
(1) Aggregation of $CCP_{\textit{word}}$ for obtaining $CCP_{\textit{claim}}$. Besides the product of probabilities, we also tried the normalized product, minimum and average probability. All these approaches perform slightly worse than the product (see Table 6).
(2) NLI context. We analyze what context is sufficient for NLI in CCP (Table 7). In addition to the standard variant in CCP, where we keep the claim that precedes the word in question, we experiment with a single target word without context and the whole sentence that precedes the target word. All variants demonstrate lower performance. No context leads to a drop in ROC-AUC of 0.02, and longer contexts by more than 0.07.
@@ 7 Conclusion
We presented a novel approach to fact-checking and hallucination detection based on token-level uncertainty quantification. According to human evaluation, our method is competitive with FactScore, a fact-checking tool that leverages an external knowledge source: we achieve similar or better results with access to only LLM outputs.
Our proposed token-level UQ method, claim conditioned probability, outperforms a number of baselines. In this method, we post-process the word distribution to mitigate the impact of uncertainty related to the variability of surface forms and uncertainty about what claim type to generate on the current step. In the constructed benchmark, where we detect hallucinations in biographies, CCP outperforms other methods for six LLMs, including GPT-3.5-turbo and GPT-4, and three languages.
@@
We strove to conduct our research according to the best established experimental and methodological practices, yet we identify several potential limitations of this work below.
@@
Third, part of our experimental results rely on a human evaluation, which may be subjective. We tried to mitigate this by creating detailed instructions for the annotators, yet a larger scale study with larger overlap may be desirable.
Fourth, our uncertainty quantification is based on tokens, while taking into account also larger units as noun or verb phrases as basic units of analysis may be better motivated linguistically.
@@
We would like to warn that our method for fact-checking is based on uncertainty quantification, and thus it is not bullet-proof, as it only reflects the internal belief of the LLM. Thus, it is of limited utility when it comes to claims that are beyond the time cutoff of the model. Moreover, the LLM might be trained on factually false data, which is beyond our control. It could be also tricked by variations in the prompt.
We also warn that our solution does not eliminate hallucinations; instead, it can be used to highlight risky parts of a text, for humans to take into account. Our intended use is towards raising awareness and promoting human-machine collaboration.
We warn that our method can be misused to unfairly moderate content. Thus, we ask researchers and potential users to exercise caution.
Appendix A Example of CCP Calculation
Figure 4 shows the example of model generation and NLI classes on the sentence from dataset sample. In this example, CCP estimates the Claim Uncertainty, not taking into account Claim order and Surface form uncertainties, in result producing more reasonable uncertainty estimate than Maximum Probability does.
Figure 4: Example part of the bibliography generated by Vicuna 13b model. The words from greedy-generated sentence are presented sequentially on the top, each non-functional word is supplemented with its alternatives and autoregressive generation probabilities. Words with probability less than 0.1% are omitted.
@@ Appendix A Example of CCP Calculation
Appendix B Fact-Checking Pipeline
Figure 5: Fact-Checking pipeline based on token-level uncertainty quantification.
@@ C.2 Datasets and Statistics
For Arabic, we generate 100 random biographies from the list of the most visited websites in Arabic Wikipedia to experiment with GPT-4. The used Arabic prompt is the translation of: “Tell me the biography of {person name}”. Then we call GPT-4 to extract atomic claims from the generated Arabic biographies. The used prompt is: “Convert the following biography into Arabic atomic factual claims that can be verified, one claim per line. Biography is: {biography}”. Arabic biographies and claims are translated into English using Google Translate. It’s worth mentioning that almost one-third of the names in the list of person names are foreign, e.g. Donald Trump, Messi, Isaac Newton, Pope Benedict XVI, etc. This can be useful to study the factuality of claims about the same persons in different Wikipedia languages. On average, GPT-4 generates 20 claims from each biography, and random two claims from each biography are verified manually (total = 200 claims).
For Jais 13b experiments, we use the same prompts used in GPT-4. We notice that the biographies generated by Jais 13b are much shorter than the ones generated by GPT-4 (almost half length). We use GPT-4 to extract claims from the generated biographies. On average, biographies generated by Jais 13b have nine claims. Out of the 100 person names, Jais 13b generates empty biographies for seven names with response messages like: “I am sorry! I cannot provide information about {name}”, or “What do you want to know exactly?”. Two random claims from each biography are verified manually (total = 186 claims).
For Chinese, we first prompt ChatGPT to generate a list of 100 famous people. Then use the same way as we have done in Arabic, but change the prompt to Chinese, to generate bios and claims. Matching of Chinese and Arabic claims is performed by aligning the matched symbols to the biography generation greedily. As FactScore pipeline only works with English claims, both Chinese and Arabic datasets were annotated manually. Only claims containing factual information were manually annotated. The statistics for this dataset are presented in Table 5
@@
Table 4: The statistics of the datasets generated out of 100 biographies for all tested LLMs for English.
@@
Table 5: The statistics of the datasets generated out of 100 biographies for all tested LLMs.
Figure 6: Example of the algorithm operation of the fact-checking method using uncertainty estimation.
Appendix D Ablation Studies
In this section, we study the influence of different components on the results of CCP method and other baselines.
D.1 Aggregation
Method ROC-AUC
$CCP_{\textit{prod}}$ 0.66 $\pm$ 0.03
$CCP_{\textit{len}}$ 0.65 $\pm$ 0.03
$CCP_{\textit{min}}$ 0.64 $\pm$ 0.03
$CCP_{\textit{mean}}$ 0.65 $\pm$ 0.03
Table 6: ROC-AUC on Vicuna 13b generation, for different normalizations.
$CCP_{\textit{prod}}(F)$ $=-\prod_{j=1}^{k}CCP_{\textit{word}}(x_{j}),$ (8)
$CCP_{\textit{len}}(F)$ $=-\exp\left(\frac{1}{k}\sum_{j=1}^{k}\log CCP_{\textit{word}}(x_{j})\right),$
$CCP_{\textit{min}}(F)$ $=-\min_{j=1}^{k}CCP_{\textit{word}}(x_{j}),$
$CCP_{\textit{mean}}(F)$ $=-\frac{1}{k}\sum_{j=1}^{k}CCP_{\textit{word}}(x_{j})$
$CCP_{\textit{prod}}$ shows better results and is the one chosen for final $CCP$ formula.
D.2 NLI Context
Method ROC-AUC
$CCP_{\textit{no context}}$ 0.64 $\pm$ 0.03
$CCP_{\textit{sent pref}}$ 0.59 $\pm$ 0.03
$CCP_{\textit{claim pref}}$ 0.66 $\pm$ 0.03
Table 7: ROC-AUC on Vicuna 13b generation, for different contexts to input words with to an NLI model.
  1. 1.
  2. 2.
  3. 3.
The results for sentence prefix provides too wide context to the NLI model, which results in NLI model focusing on the the contexts instead of the greedy words and its alternatives, thus outputting lots of entailment classes. On the other hand, the method without context does not provide enough context to calculate NLI class with more quality. $CCP_{\textit{claim pref}}$ shows better results and is the one chosen for final $CCP$ formula.
D.3 Functional Words Handling
Method ROC-AUC
$CCP_{\textit{confident}}$ 0.66 $\pm$ 0.03
$CCP_{\textit{ignore}}$ 0.63 $\pm$ 0.03
$MP_{\textit{confident}}$ 0.59 $\pm$ 0.05
$MP_{\textit{ignore}}$ 0.60 $\pm$ 0.05
Table 8: ROC-AUC on Vicuna 13b generation, for methods of handling functional words.
  1. 1.
  2. 2.
D.4 The Number of Alternatives
Figure 7: ROC-AUC between FactScore classes and claim-level CCP method, dependent on the number of sampled words $n$ to consider.
Appendix E Additional Experimental Results
Here we show additional results related to the performance of our pipeline and CCP method. In the Figure 8 we show the dependence of the percentage of supported claims in the generated LLM response (in this case for Vicuna 13b) as a function of its length. As we can see, the model is more often wrong when the generation length increases, which may be due to the generation of additional facts.
In the Figure 9, we show the performance quality of our CCP method on Chinese data on the Yi 6b model. We see that, as for other English models, our method outperforms the alternatives over the entire generation length in terms of ROC-AUC.
Appendix F FactScore Annotation System

### v2

@@ 5 Benchmark for Evaluation of Claim-Level UQ Methods
We evaluate claim-level UQ techniques and their ability to spot hallucinations on the task of generating biographies. In relevant previous work Manakul et al. (2023), the authors generate biographies with one LLM (GPT-3), manually annotate sentences for factuality, and quantify uncertainty of a different “proxy” model. Factuality labels are then used to evaluate the quality of uncertainty scores. We argue that such an approach based on a proxy model introduces a big discrepancy between the generated text and what a proxy LLM actually wants to generate, which results in biased UQ evaluation results. To make the evaluation as close as possible to the real-world scenario, we allow unrestricted generation of biographies from LLMs.
Unrestricted generation complicates automatic evaluation of the fact-checking pipeline, because obtaining gold standard annotation requires manual annotation of all outputs from each model. Therefore, in addition to manual annotation, we annotate claims in generated texts automatically using FactScore – a fact-checking tool Min et al. (2023), which has access to an external knowledge source. Using FactScore, enables completely automatic evaluation and allows us to scale up experiments.
We generate LLM responses in English, Chinese, Arabic, and Russian to $100$ biography prompts. The typical biography prompt is Give me a biography for <person name> in different languages. The set of people was generated by asking GPT-4 to list the most famous people since $1900$. The maximum generation length is set to $256$ tokens. If the last sentence of the generation is unfinished (i.e. does not end with any punctuation), it is discarded. We generate responses for the following LLMs: (for English) Vicuna 13b Zheng et al. (2023), Mistral 7b Jiang et al. (2023), Jais 13b Sengupta et al. (2023), and GPT-3.5-turbo Ouyang et al. (2022); (for Chinese) Yi 6b Young et al. (2024); (for Arabic) Jais 13b and GPT-4; (for Russian) Vikhr-instruct-0.2 7b Nikolich et al. (2024).
@@
Model Mistral 7b Vicuna 13b Jais 13b GPT-3.5-turbo
CCP (ours) 0.34 $\pm$ 0.05 0.24 $\pm$ 0.04 0.33 $\pm$ 0.07 0.14 $\pm$ 0.01
Maximum Prob. 0.26 $\pm$ 0.04 0.18 $\pm$ 0.05 0.24 $\pm$ 0.05 0.13 $\pm$ 0.02
Perplexity 0.27 $\pm$ 0.03 0.17 $\pm$ 0.04 0.21 $\pm$ 0.05 0.11 $\pm$ 0.03
Token Entropy 0.30 $\pm$ 0.06 0.18 $\pm$ 0.04 0.24 $\pm$ 0.06 0.12 $\pm$ 0.02
P(True) 0.24 $\pm$ 0.03 0.21 $\pm$ 0.06 0.19 $\pm$ 0.06 0.17 $\pm$ 0.02
Table 2: PR-AUC (considering the Not Supported class as positive) of claim-level UQ methods with FactScore labels as the ground truth (English).
@@
We decompose the generated text into atomic claims using GPT-4. For each claim, we map all its words back to generated text to access the corresponding token logits. Not all claims perfectly match to the original response. For example, for Vicuna 13b around $5\%$ of all claims do not successfully match because ChatGPT abstained to respond or outputted words not present in the original. We consider only successfully matched claims.
For English, we are able to perform annotation completely automatically: each atomic claim is classified by FactScore as supported or not supported. The underlying FactScore model is “retrieval+ChatGPT” with a dump of Wikipedia articles as an external knowledge source. We also manually annotated $100$ claims from English biographies produced by the Vicuna 13b model, $183$ claims from Arabic biographies, $146$ claims from Russian biographies, and $1603$ claims in Chinese. Each of the statements was checked by two annotators with access to the corresponding Wikipedia article. The final label is set to “supported” only if both annotators label it as so.
The statistics of the resulting datasets with automatically-labeled claims in English are presented in Table 7, and the statistics of manually annotated datasets are presented in Table˜7. The majority of claims in the model output are correct, with 662929% of hallucinations. The automatic pipeline for evaluation of UQ methods using FactScore is illustrated in Figure 6 in Appendix C.
@@ 6.1 Experimental Setup
Fact-checking of atomic claims is framed as a binary classification task, where uncertainty scores serve as predictors of non-factuality, and FactScore or human labels serve as ground truth. The evaluation metric is ROC-AUC and PR-AUC (unsupported claims as a positive class).
For the CCP method, NLI scores are calculated using the DeBERTa-large model He et al. (2021) fine-tuned for this task [^2] https://huggingface.co/microsoft/deberta-large-mnli. The number of alternatives used in CCP is $K=10$, except for GPT-3.5-turbo and GPT-4, as the OpenAI API does not allow us to retrieve more than $5$ alternatives from the token distribution. Details on hardware and computational resources spent for experiments can be found in Appendix F.
@@ 6.2 Results for English on the FactScore Annotation
The main results of the experiments with FactScore labels for English are presented in Tables 1 and 2. The proposed CCP method outperforms all other UQ techniques for each of the considered LLMs, with only exception in the PR-AUC metric for the GPT-3.5-turbo model, where the P(True) approach exhibits the highest performance. The underperformance of CCP for GPT-3.5-turbo may be attributed to the limited number of token options and their associated logits available through the OpenAI API. The best overall improvement from CPP is obtained for Jais 13b, where it outperforms the closest competitor by 0.07 ROC-AUC and 0.09 PR-AUC.
We further analyze the performance by plotting ROC-AUC as a function of the number of considered sentences from the beginning of the generated text (see Figure 3). We note that the quality of each method decreases as the number of considered sentences increases. This may be due to the fact that the model tends to start the response with easy-to-know and hence reliable claims and as it generates more text, it has to produce more complex and less reliable statements, which is illustrated for the Vicuna 13b model in Figure 8 in the appendix. For the majority of cases, CCP outperforms other methods, except when we consider only the first two and the first five sentences generated by GPT-3.5-turbo.
@@ 6.3 Multilingual Results on Manual Annotation
Multilingual results based on manual annotation are presented in Tables˜3 and 4 and in Figure˜9. Using manual annotation for English, we can also evaluate the performance of FactScore itself. The accuracy of automatic annotation is $77.2\%$ and ROC-AUC is $0.72$. A detailed analysis of FactScore mistakes is presented in Appendix C.3.
For human annotation, the performance of all UQ methods appears to be even slightly higher than for the labels obtained using FactScore (Table 3). Moreover, CCP outperforms FactScore itself by $0.06$ ROC-AUC. These results demonstrate that in the task of detecting LLM hallucinations, UQ techniques can be a strong alternative to fact-checking tools with an external knowledge source.
For Chinese, Arabic, and Russian, CCP also outperforms the baselines. For the Chinese Yi 6b model, from Figure 9, we can see that the gap between CCP and the baselines is especially significant for the several first claims. When considering more claims, CCP still clearly outperforms the Maximum Probability baseline, but the P(True) baseline substantially reduces the gap. For Arabic and Jais, CCP outperforms the closest competitor by $0.05$ ROC-AUC. On GPT-4 generations for Arabic, the metrics for all methods are low. We explain this observation by a small ratio of non-factual claims in the GPT-4 output. For Russian generations with the Vikhr model, CCP confidently outperforms Maximum Probability, which is its closes competitor, by $0.05$ ROC-AUC.
@@
Model Yi 6b, Chinese Jais 13b, Arabic GPT-4, Arabic Vikhr 7b, Russian CCP (ours) 0.64 $\pm$ 0.03 0.66 $\pm$ 0.02 0.56 $\pm$ 0.05 0.68 $\pm$ 0.04 Maximum Prob. 0.52 $\pm$ 0.03 0.59 $\pm$ 0.02 0.55 $\pm$ 0.08 0.63 $\pm$ 0.04 Perplexity 0.51 $\pm$ 0.04 0.56 $\pm$ 0.02 0.54 $\pm$ 0.08 0.58 $\pm$ 0.04 Token Entropy 0.57 $\pm$ 0.05 0.61 $\pm$ 0.02 0.48 $\pm$ 0.06 0.55 $\pm$ 0.03 P(True) 0.63 $\pm$ 0.04 0.61 $\pm$ 0.02 0.50 $\pm$ 0.06 0.58 $\pm$ 0.03
@@ 6.4 Ablation Studies
(1) Aggregation of $CCP_{\textit{word}}$ for obtaining $CCP_{\textit{claim}}$. Besides the product of probabilities, we also tried the normalized product, minimum, and average probability. All these approaches perform slightly worse than the product (see Table 9).
(2) NLI model. We investigate the influence of the specific NLI model on the performance of CCP. Table 10 shows that CCP’s effectiveness is not critically dependent on the complexity of the NLI model employed. Notably, even a relatively small model with 22M parameters maintains strong performance without any degradation.
(3) NLI context. We analyze what context is sufficient for NLI in CCP (Table 11). In addition to the standard variant in CCP, where we keep the claim that precedes the word in question, we experiment with a single target word without context and the whole sentence that precedes the target word. All variants demonstrate lower performance. No context results in a drop of $0.02$ ROC-AUC, and longer contexts – of more than $0.07$.
@@ 6.5 Qualitative Analysis
6.6 Computational Efficiency
Method NLI model Runtime
parameters
MPMP 18.5 $\pm$ 0.8 sec
CCPCCP 350M 20.1 $\pm$ 0.9 sec
CCPCCP 22M 19.1 $\pm$ 0.8 sec
Table 5: The runtime of Maximum Probability and CCP methods on 100 biographies in English.
To demonstrate the computational efficiency of CCP, we compare it to the fastest UQ method – Maximum Probability. Experiments were conducted using a dataset of 100 biographies and Mistral 7b Jiang et al. (2023). To ensure a fair comparison, we focus solely on the runtime of generating biographies and calculating the respective uncertainty scores for each claim. Time spent on claim extraction and matching is excluded. The experiments utilized two 32GB V100 GPUs. Each biography was processed in a single batch.
MP does not introduce notable overhead over the generation process, as it only aggregates produced logits. CCP involves running an NLI model for each of 10 token candidates per token position, which introduces some overhead.
Table 5 presents the runtime comparison. We see that NLI does not make a substantial impact. Using the default microsoft/deberta-large-mnli model (350M parameters) results in 8% increase of the runtime compared to MP. Using the smaller cross-encoder/nli-deberta-v3-xsmall model (22M parameters), which achieves comparable UQ performance, reduces the computational overhead to only 3%.
@@ 7 Conclusion
We presented a novel approach to fact-checking and hallucination detection based on token-level uncertainty quantification. According to human evaluation, our approach is competitive with FactScore, a fact-checking tool that leverages an external knowledge source: we achieve similar or better results with access to only LLM outputs.
We proposed a computationally efficient token-level and claim-level UQ method, Claim Conditioned Probability that outperforms a number of baselines in fact-checking. In this method, we post-process the word distribution to mitigate the impact of uncertainty related to the variability of surface forms and uncertainty about what claim type to generate on the current step. In the constructed benchmark, where we detect hallucinations in biographies, CCP outperforms other methods for seven LLMs, including GPT-3.5-turbo and GPT-4, and four languages. We also demonstrate that computational overhead of CCP might be as low as 3% of the LLM inference runtime.
@@
While this work has been conducted according to experimental and methodological best practice, there are several potential limitations.
@@
Third, part of our experimental results rely on a human evaluation, which may be subjective. We tried to mitigate this by creating detailed instructions for the annotators, but a larger-scale study with larger overlap would further strengthen the results.
Fourth, our uncertainty quantification is based on tokens, where taking into account also larger units such as noun or verb phrases as basic units of analysis may be better motivated linguistically.
Fifth, in this work, we do not consider the calibration of CCP scores. We note that calibration on its own does not provide the information about the performance in the practical task that we are interested in – fact-checking. At the same time, CCP could be post-calibrated in the same way as any other probabilities or their surrogates.
@@
We would like to caution that our method for fact-checking is based on uncertainty quantification, and thus it is not bullet-proof, as it only reflects the internal model state of the LLM. Thus, it is of limited utility when it comes to claims that are beyond the time cutoff of the model. Moreover, the LLM might be trained on factually false data, which is beyond our control. It could be also tricked by variations in the prompt.
We also caution that our solution does not eliminate hallucinations; instead, it can be used to highlight risky parts of a text, for humans to take into account. Our intended use is towards raising awareness and promoting human-machine collaboration.
Finally, our method can be misused to unfairly moderate content. Thus, we ask researchers and potential users to exercise due caution.
Appendix A Fact-Checking Pipeline
Figure 4: The scheme of the fact-checking pipeline based on UQ.
Appendix B Example of CCP Calculation
Figure 5 demonstrates an example of the LLM generation process and CCP calculation process. CCP quantifies the Claim Uncertainty, not taking into account the Claim order and Surface form uncertainties. As a result, CCP produces better uncertainty scores than Maximum Probability.
Figure 5: Example of the Vicuna 13b generation process and CCP calculation process part. The words from the greedy-generated sentence are presented sequentially on the top, each non-functional word is supplemented with its alternatives and autoregressive generation probabilities. Words with probability less than 0.1% are omitted.
@@ C.1 Benchmark Construction Pipeline
Figure 6: A visualization of the benchmark construction pipeline based on FactScore with an example.
@@ C.2 Datasets and Statistics
Table 6: The statistics of the datasets generated from 100 biographies using all tested English LLMs and annotated automatically.
@@
Table 7: The statistics of the datasets generated from 100 biographies using all tested English LLMs and annotated manually.
Since FactScore only supports English, for Arabic, Chinese, and Russian, we generate biographies of well-known people and annotate them only manually. We also manually annotate 100 English claims generated by Vicuna 13b. The statistics for the annotated datasets are presented in Table 7.
For Arabic, using GPT-4, we generate 100 biographies of people randomly selected from the list of the most visited websites in Arabic Wikipedia. The used Arabic prompt is the translation of: “Tell me the biography of {person name}”. To extract claims, we prompt GPT-4 in the following way: “Convert the following biography into Arabic atomic factual claims that can be verified, one claim per line. Biography is: {biography}”. Arabic biographies and claims are translated into English using Google Translate. It is worth mentioning that almost one-third of the names in the list of person names are foreign, e.g. Donald Trump, Messi, Isaac Newton, Pope Benedict XVI, etc. On average, GPT-4 generates 20 claims from each biography, and random two claims from each biography are verified manually (total = 200 claims).
For Jais 13b experiments, we use the same prompts used for GPT-4. We notice that the biographies generated by Jais 13b are much shorter than the ones generated by GPT-4 (almost half-length). Similarly, we use GPT-4 to extract claims from the generated biographies. On average, biographies generated by Jais 13b have nine claims. Jais 13b generates empty biographies for seven names (out of 100) with response messages like: “I am sorry! I cannot provide information about {name}”, or “What do you want to know exactly?”. Two random claims from each biography are verified manually (total = 186 claims).
For Chinese, we first prompt ChatGPT to generate a list of 100 famous people. Then use the same way as we have done in Arabic, but change the prompt to Chinese, to generate biographies and claims. We use Yi 6b to generate texts and GPT-4 to split them into atomic claims.
For Russian, we conduct a similar approach, prompting ChatGPT to generate a list of 100 famous people and checking the result to obtain representative personalities from different areas such as science, sport, literature, art, government activity, cinematography, heroes, etc. A balanced list of famous people in different professional categories was obtained. For these people, we generate biographies using the Vikhr 7b model Nikolich et al. (2024).
C.3 FactScore Annotation